1. A rice breeder obtained a triple heterozygote carrying the three recessive alleles for albino flowers (al), brown awns (b), and fuzzy leaves (fu), all paired with their normal wild -type alleles. This triple heterozygote was testcrosses .The progeny phenotypes were:


Number Phenotypes Genotype

170 wild type al+ b+ fu+


150 albino, brown, al b fu

fuzzy


5 brown al+ b fu+


3 albino, fuzzy al b+ fu


710 albino al b+ fu+


698 brown, fuzzy al+ b fu


42 albino, brown al+ b+ fu


38 albino, brown al b fu+


I. Are any of the genes linked?


II. If they are linked, draw a map labelled with map distances

III.The triple heterozygote was originally made by crossing two pure lines . What were their genotypes?

IV. Calculate interference and say what you think of its signicane.




1
Expert's answer
2022-01-31T05:33:01-0500

Step1

a) Yes based on the data, the three gene are linked. The most common Class of progeny represent the " parental" chromosome( al b^+ fu^+ and al^+ bfu) while the least common class represent the double cross event( al^+ bfu^+ ad alb^+ fu) can be used to deduce the gene order.

In this comparison, we can say that the middle gene is fu to calculate the map distances we should first determine the various recombinant classes; This is shown below




Steps 11

Based on this data, there are 170+150+5+3=328 recombinants between al and fu with a map distance of ( 328/total progeny)×100%=18M.U on the other hand there are 48+38+5+3=88 recombinants between fu and b with a map distance of 4.8M.U



Stepiii) The parentals,or two pure lines are indicated by the most common classses of progeny. In this case the cross was between: al fu^+ b^+/ al fu^+ b^+ × al^+ fub/ al^+ fub

Stepiv) Test cross between a triple heterozygote suggests that the three genes are linked